GNU Accounting Utilities

Version 6.3.5 26 May 1998

Noel Cragg (noel@gnu.ai.mit.edu)

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1993, 1996, 1997 Free Software Foundation, Inc. Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies. Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one. Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that this permission notice may be stated in a translation approved by the Foundation.

Preface

Way back a long time ago, Thompson and Ritchie were sitting opposite one another at the commissary, sipping coffees and discussing their evolving behemoth.

“This behemoth of ours,” said Ken, “is becoming rather popular, wouldn’t you say?” “Yes,” said Dennis. “Every time I want to do a compilation, I have to wait for hours and hours. It’s infuriating.” They both agreed that the load on their system was too great. Both sighed, picked up their mugs, and went back to the workbench. Little did they know that an upper-management type was sitting just within earshot of their conversation.

“We are AT&T Bell Laboratories, aren’t we?” the upper-management type thought to himself. “Well, what is our organization best known for?” The brill-cream in his hair glistened. “Screwing people out of lots of money, of course! If there were some way that we could keep tabs on users and charge them through the nose for their CPU time...”

The accounting utilities were born.

Seriously though, the accouting utilities can provide a system administrator with useful information about system usage—connections, programs executed, and utilization of system resources.

Information about users—their connect time, location, programs executed, and the like— is automatically recored in files by init and login. Four of them are of interest to us: wtmp, which has records for each login and logout; acct, which records each command that was run; usracct and savacct, which contain summaries of the information in acct by user and command, respectively. Each of the accounting utilities reports or summarizes information stored in these files.

ac prints statistics about users’ connect time. ac can tell you how long a particular user or group of users were connected to your system, printing totals by day or for all of the entries in the wtmp file.

accton turns accounting on or off.

last lists the logins on the system, most recent first. With last, you can search the wtmp file for a particular user or terminal name (to which the user was connected). Of special interest are two fake users, ‘reboot’ and ‘shutdown’, which are recorded when the system is shut down or reboots.

lastcomm lists the commands executed on the system, most recent first, showing the run state of each command. With last, you can search the acct file for a particular user, terminal, or command.

sa summarizes the information in the acct file into the savacct and usracct file. It also generates reports about commands, giving the number of invocations, cpu time used, average core usage, etc.

dump-acct
dump-utmp

display acct and utmp files in a human-readable format.

For more detailed information on any of these programs, check the chapter with the program title.

A Note on File Names and Locations

The wtmp and acct files seem to live in different places and have different names for every variant of u*x that exists. The name wtmp seems to be standard for the login accounting file, but the process accounting file might be acct or pacct on your system. To find the actual locations and names of these files on your system, specify the --help flag to any of the programs in this package and the information will dumped to standard output.

Regardless of the names and locations of files on your system, this manual will refer to the login accounting file as wtmp and the process accounting files as acct, savacct, and

usracct.

History of the Accounting Utilities

I don’t have any idea who originally wrote these utilities. If anybody does, please send some mail to noel@gnu.ai.mit.edu and I’ll add your information here!

Since the first alpha versions of this software in late 1993, many people have contributed to the package. They are (in alphabetical order):

Eric Backus <ericb@lsid.hp.com>

Suggested fixes for HP-UX 9.05 using /bin/cc: configure assumed you were using gcc and tacked on -Wall etc. He also noticed that file_rd.c was doing pointer arithmetic on a void * pointer (non-ANSI).

Christoph Badura <bad@flatlin.ka.sub.org>

Christoph was a BIG HELP in computing statistics, most notably k*sec stuff! He also did Xenix testing and contributed some Makefile fixes and output optimizations.

Michael Calwas <calwas@ttd.teradyne.com>

Fixed bugs in mktime.c.

Derek Clegg <dclegg@apple.com>

Suggested the simple, elegant fix for * rd never used brain-damage.

Alan Cox <iiitac@pyr.swan.ac.uk>

Original Linux kernel accounting patches.

Scott Crosby <root@hypercube.res.cmu.edu>

Suggested idea behind --sort-real-time for sa.

Solar Designer <solar@false.com>

Added code for --ahz flag in lastcomm and sa.

Dirk Eddelbuettel <edd@miles.econ.queensu.ca>

Managed bug-fixes & etc. for Debian distribution, as well as the architect of merge of GNU + Debian distributions. A big thanks to Dirk for kicking me back into gear again after a long period of no work on this project.

Jason Grant <jamalcol@pc-5530.bc.rogers.wave.ca>

Identified a buffer-overrun bug in sa.

Kaveh R. Ghazi <ghazi@caip.rutgers.edu>

Tested the package on many systems with compilers other than gcc. Fixed K&R C support.

Susan Kleinmann <sgk@sgk.tiac.net>

Contributed excellent man pages!

Alexander Kourakos <Alexander@Kourakos.com>

Inspired the --wide option for last.

Marek Michalkiewicz <marekm@i17linuxb.ists.pwr.wroc.pl>

Suggested the --ip-address flag for last.

David S. Miller <davem@caip.rutgers.edu>

Noticed missing GNU-standard makefile rules.

Walter Mueller <walt@pi4.informatik.uni-mannheim.de>

Noticed install target was missing, and corrected a typo for prefix in Makefile.in.

Ian Murdock <imurdock@gnu.ai.mit.edu>

Tracked down miscellaneous bugs in sa.c under Linux. Added Debian package maintenance files.

Tuomo Pyhala <tuomo@lesti.kpnet.fi>
          Reported buggy --strict-match flag in lastcomm.
Luc I. Suryo <root@patriots.nl.mugnet.org>

Suggested the --user flag for lastcomm.

Pedro A M Vazquez <vazquez@iqm.unicamp.br>

Fixed bugs in sa.c and tested under FreeBSD.

Marco van Wieringen <Marco.van.Wieringen@mcs.nl.mugnet.org>

Modified (wrote?) Linux kernel accounting patches.

1 ac

The ac command prints out a report of connect time (in hours) based on the logins/logouts in the current wtmp file. A total is also printed out.

The accounting file wtmp is maintained by init and login. Neither of these programs creates the file; if the file is not there, no accounting is done. To begin accounting, create the file with a length of zero. NOTE: the wtmp file can get really big, really fast. You might want to trim it every once and a while.

GNU ac works nearly the same u*x ac, though it’s a little smarter in its printing out of daily totals—it actually prints every day, rather than skipping to the date of the next entry in the wtmp file.

1.1 Flags

All of the original ac’s options have been implemented, and a few have been added. Normally, when ac is invoked, the output looks like this:

             total 93867.14

where total is the number of hours of connect time for every entry in the wtmp file. The rest of the flags modify the output in one way or another.

-d
--daily-totals

Print totals for each day rather than just one big total at the end. The output looks like this:

               Jul  3  total     1.17
               Jul  4  total     2.10
               Jul  5  total     8.23
               Jul  6  total     2.10
               Jul  7  total     0.30
-p
--individual-totals

Print time totals for each user in addition to the usual everything-lumped-into-one value. It looks like:

                       bob       8.06
                       goff      0.60
                       maley     7.37
                       root      0.12
                       total    16.15

people Print out the sum total of the connect time used by all of the users included in people. Note that people is a space separated list of valid user names; wildcards are not allowed.

-f filename
--file filename

Read from the file filename instead of the system’s wtmp file.

--complain

When the wtmp file has a problem (a time-warp, missing record, or whatever), print out an appropriate error.

--reboots

Reboot records are not written at the time of a reboot, but when the system restarts; therefore, it is impossible to know exactly when the reboot occurred. Users may have been logged into the system at the time of the reboot, and many ac’s automatically count the time between the login and the reboot record against the user (even though all of that time shouldn’t be, perhaps, if the system is down for a long time, for instance). If you want to count this time, include the flag. To make ac behave like the one that was distributed with your OS, include this flag.

--supplants

Sometimes a logout record is not written for a specific terminal, so the time that the last user accrued cannot be calculated. If you want to include the time from the user’s login to the next login on the terminal (though probably incorrect), include this flag. To make ac behave like the one that was distributed with your OS, include this flag.

--timewarps

Sometimes, entries in a wtmp file will suddenly jump back into the past without a clock change record occurring. It is impossible to know how long a user was logged in when this occurs. If you want to count the time between the login and the time warp against the user, include this flag. To make ac behave like the one that was distributed with your OS, include this flag.

--compatibility

This is shorthand for typing out the three above options.

-a
--all-days

If we’re printing daily totals, print a record for every day instead of skipping intervening days where there is no login activity. Without this flag, time accrued during those intervening days gets listed under the next day where there is login activity.

-y
--print-year

Print out the year when displaying dates.

--print-zeros

If a total for any category (save the grand total) is zero, print it. The default is to suppress printing.

--debug Print verbose internal information.

--tw-leniency value

Set the time warp leniency value (in seconds). Records in wtmp files might be slightly out of order (most notably when two logins occur within a one-second period – the second one gets written first). By default, this value is set to 1 second. Some wtmp’s are really screwed up (Suns) and require a larger value here. If the program notices this problem, time is not assigned to users unless the --timewarps flag is used. See the Problems section for more information.

--tw-suspicious value

Set the time warp suspicious value (in seconds). If two records in the wtmp file are farther than this number of seconds apart, there is a problem with the wtmp file (or your machine hasn’t been used in a year). If the program notices this problem, time is not assigned to users unless the --timewarps flag is used.

-V
--version

Print ac’s version number.

-h

--help Print ac’s usage string and default locations of system files to standard output.

1.2 Problems

For no fault of ac’s, if two logins occur at the same time (within a second of each other), each login process will try to write an entry to the wtmp file. With file system overhead, it is forseeable that the entries would get written in the wrong order. GNU ac automatically compensates for this, but some other acs may not... beware.

The FTP Problem

I’ve tested the standard ac in Ultrix 4.2 (DECstation/DECsystem), SunOS 4.1.1 (Sun3, Sun4, Sparc), Mach 2.5 (Omron/Luna), and DomainOS 10.3 (DN3500). All of these acs have trouble parsing entries in which the line is ftpxxxx (xxxx being some number). Whenever these acs see one of these entries, they log everyone out at the time of the entry. HOW IT HAPPENS: if there is a user logged into the machine when an ftp connection occurs, (minimally) you’ll get a login record for the user, a login record for the ftp connection, and the logouts for both afterwards (in either order). TANGIBLE RESULT: the user who was logged in gets ’logged out’ at the time the ftp connection begins, and none of the time spent during or after the ftp connection. Therefore, when you run GNU ac, the totals will most likely be greater than those of your system’s ac (provided you specify the other flags that will make GNU ac behave like the system’s).

The Shutdown/Reboot Problem

On Suns, init is a little screwed up. For some reason, after a shutdown record is written, a reboot record is written with a time-stamp before the shutdown (less than 30 seconds, usually). TANGIBLE RESULT: GNU ac will notice the problem, log everyone out (you can specify if you want the time to be added to the user’s total) and begin a new day entry based on the time of the out-of-sync record. If you try to print out daily totals, you’ll notice that some days might have two or more entries. SOLUTION: To fix this, a timewarp leniency value has been implemented. If any record is out of order by this number of seconds (defaults to 60) it gets ignored. If you need to change this value (if you think the totals are off because the value is too high), you can change it using the ‘--timewarp-value’ flag. The rationale for the 60 second default is that of all of the machines with this problem, the largest timewarp was 45.

Stupid System V Machines

Some ac’s on System V machines (I’ve tried SGI Indigo & SGI Indy) forget to pay attention to the ut_type field in a struct utmp. As such, they chalk up a lot of time to non-existant processes called LOGIN or runlevel. TANGIBLE RESULT: The amount of total time reported by the system’s ac is really off. Often, it’s several times greater than what it should be. SOLUTION: GNU ac always pays attention to the ut_type record, so there’s no possibility of chalking up time to anything but user processes.

2 accton

accton turns process accounting on or off. To save process accounting information in accountingfile, use:

   accton accountingfile

If called with no arguments, it will, by default, stop process accounting.

2.1 Flags

-V
--version

Print accton’s version number.

-h

--help Print accton’s usage string and default locations of system files to standard output.

3 last

last looks through the wtmp file (which records all logins/logouts) and prints information about connect times of users. Records are printed from most recent to least recent. Records can be specified by tty and username. tty names can be abbreviated: ‘last 0’ is equivalent

to ‘last tty0’.

Multiple arguments can be specified: ‘last root console’ will print all of the entries for the user root and all entries logged in on the console tty.

The special users reboot and shutdown log in when the system reboots or (surprise) shuts down. ‘last reboot’ will produce a record of reboot times.

If last is interrupted by a quit signal, it prints out how far its search in the wtmp file had reached and then quits:

     weerapan  ttyq6    132.162.32.37    Mon Feb 15 19:07 - 19:21  (00:13)
     weerapan  ttyq6    132.162.32.37    Mon Feb 15 19:07 - 19:21  (00:13)
     interrupted at Mon Feb 15 19:07:52 1993

3.1 Flags

This program implements the features of regular u*x last with a few extra flags. When last is invoked with no arguments, the output looks like this:

     gr151     ttyp2    ray.cs.oberlin.e Tue Feb 16 17:40   still logged in
     jhoggard  ttyp2    csts.cs.oberlin. Tue Feb 16 17:39 - 17:39  (00:00)
     jstarr    ttyp1    UNIX5.ANDREW.CMU Tue Feb 16 17:38   still logged in
     jberman   ttypb    132.162.32.25    Tue Feb 16 17:34   still logged in
     alee      ttyp7    csts.cs.oberlin. Tue Feb 16 17:34   still logged in
     jbrick    ttyp2    ocvaxa.cc.oberli Tue Feb 16 17:33 - 17:36  (00:03)
     mbastedo  ttypc    ocvaxa.cc.oberli Tue Feb 16 17:25 - 17:26  (00:01)
     rgoodste  ttypb    ocvaxa.cc.oberli Tue Feb 16 17:22 - 17:26  (00:03)
     huttar    ttyp9    lobby.ti.com     Tue Feb 16 17:19   still logged in
     klutz     ttyp3    132.162.32.25    Tue Feb 16 17:14   still logged in
--no-truncate-ftp-entries

When printing out the information, don’t chop the number part off of ftpxxxx entries.

-number
-n number
--lines number

Limit the number of lines that last prints.

-f filename
--file filename

Read from the file filename instead of the system’s wtmp file.

-y
--print-year

Print out the year when displaying dates.

-s
--print-seconds

Print out seconds when displaying dates and durations.

--complain

When the wtmp file has a problem (a time-warp, missing record, or whatever), print out an appropriate error.

-x
--more-records

Print out run level changes, shutdowns, and time changes in addition to the normal records.

-a
--all-records

Print out all records in the wtmp file.

-i
--ip-address

Some machines store the IP address of a connection in a utmp record. Enabling this option makes last print the IP address instead of the hostname.

--tw-leniency value

Set the time warp leniency value (in seconds). See the ac chapter for information.

--tw-suspicious value

Set the time warp suspicious value (in seconds). See the ac chapter for information.

-w

--wide By default, last tries to print each entry within in 80 columns. Use this option to instruct last to print out the fields in the wtmp file with full field widths.

--debug Print verbose internal information.

-V
--version

Print last’s version number.

-h

--help Print last’s usage string and default locations of system files to standard output.

3.2 Problems

The Clock Change Problem

Of the lasts I’ve tried, all of them have had problems parsing a system clock change. Instead of modifying the entries that have been read, they just ignore the change and give you incorrect values. GNU last knows about clock changes and prints the correct times. TANGIBLE RESULT: if you diff the output of your last and GNU last, entries after (before, rather) a clock change will be off by the amount of the clock change.

The Ftp Problem

Most lasts that I’ve examined have the same problem here as ac does—they log everyone out as soon as they see an ftp entry. TANGIBLE RESULT: GNU last will reflect the correct time spent in an ftp session, so the totals that it gives will most likely be greater than those given by the system last.

4 lastcomm

lastcomm prints out information about previously executed commands. If no arguments are specified, lastcomm will print info about all of the commands in the acct file (the record file). If called with a command name, user name, or tty name, only records containing those items will be displayed. For example, to find out which users used command ‘a.out’ and which users were logged into ‘tty0’, type:

   lastcomm a.out tty0

This will print any entry for which ‘a.out’ or ‘tty0’ matches in any of the record’s fields (command, name, or tty). If you want to find only items that match ALL of the arguments on the command line, you must use the ’–strict-match’ option. For example, to list all of the executions of command ‘a.out’ by user ‘root’ on terminal ‘tty0’, type:

   lastcomm --strict-match a.out root tty0

The order of the arguments is not important.

For each entry the following information is printed:

• command name of the process

• flags, as recorded by the system accounting routines: − S command executed by super-user − F command executed after a fork but without a following exec − C command run in PDP-11 compatibility mode (VAX only) − D command terminated with the generation of a core file − X command was terminated with the signal SIGTERM

• the name of the user who ran the process

• time the process exited

4.1 Flags

This program implements the features of regular u*x lastcomm with a few extra flags. When lastcomm is invoked without arguments, the output looks like this:

     nslookup         jberman  ttypb      0.03 secs Tue Feb 16 19:23
     comsat           root     __0.03 secs Tue Feb 16 19:19
     uptime           ctilburg __0.11 secs Tue Feb 16 19:23
     sh          F    ctilburg __0.02 secs Tue Feb 16 19:23
     sleep            ctilburg __0.02 secs Tue Feb 16 19:22
     ls               noel     ttyp4      0.19 secs Tue Feb 16 19:23
--strict-match

Print only entries that match all of the arguments on the command line.

--user name

List records for user with name. This is useful if you’re trying to match a username that happens to be the same as a command (e.g., ed).

--command name

List records for command name.

--tty name

List records for tty name.

-f filename
--file filename

Read from the file filename instead of the system’s acct file.

--ahz hz Use this flag to tell the program what AHZ should be (in hertz). This option is useful if you are trying to view an acct file created on another machine which has the same byte order and file format as your current machine, but has a different value for AHZ.

--debug Print verbose internal information.

--version

Print lastcomm’s version number.

--help Print lastcomm’s usage string and default locations of system files to standard output.

5 sa

sa summarizes information about previously executed commands as recorded in the acct file. In addition, it condenses this data into the savacct summary file, which contains the number of times the command was called and the system resources used. The information can also be summarized on a per-user basis; sa will save this information into usracct. Usage:

   sa [opts] [file]

If no arguments are specified, sa will print information about all of the commands in the acct file. If command names have unprintable characters, or are only called once, sa will sort them into a group called ***other.

If called with a file name as the last argument, sa will use that file instead of acct.

By default, sa will sort the output by sum of user and system time.

The output fields are labeled as follows:

cpu sum of system and user time in cpu seconds

re “real time” in cpu seconds

k cpu-time averaged core usage, in 1k units

avio average number of I/O operations per execution

tio total number of I/O operations

k*sec cpu storage integral (kilo-core seconds)

u user cpu time in cpu seconds

s system time in cpu seconds

Note that these column titles do not appear in the first row of the table, but after each numeric entry (as units of measurement) in every row. For example, you might see 79.29re, meaning 79.29 cpu seconds of “real time.”

An asterisk will appear after the name of commands that forked but didn’t call exec.

5.1 Flags

The availability of these program options depends on your operating system. In specific, the members that appear in the struct acct of your system’s process accounting header file (usually acct.h) determine which flags will be present. For example, if your system’s struct acct doesn’t have the ac_mem field, the installed version of sa will not support the

--sort-cpu-avmem, --sort-ksec, -k, or -K options.

In short, all of these flags may not be available on your machine.

-a
--list-all-names

Force sa not to sort those command names with unprintable characters and those used only once into the ‘***other’ group.

-b
--sort-sys-user-div-calls

Sort the output by the sum of user and system time divided by the number of calls.

-c
--percentages

Print percentages of total time for the command’s user, system, and real time values.

-d
--sort-avio

Sort the output by the average number of disk I/O operations.

-D
--sort-tio

Print and sort the output by the total number of disk I/O operations.

-f
--not-interactive

When using the --threshold option, assume that all answers to interactive queries will be affirmative.

-i
--dont-read-summary-file

Don’t read the information in savacct.

-j
--print-seconds

Instead of printing total minutes for each category, print seconds per call.

-k
--sort-cpu-avmem

Sort the output by cpu time average memory usage.

-K
--sort-ksec

Print and sort the output by the cpu-storage integral.

-l
--separate-times

Print separate columns for system and user time; usually the two are added together and listed as cpu.

-m
--user-summary

Print the number of processes and number of CPU minutes on a per-user basis.

-n
--sort-num-calls

Sort the output by the number of calls. This is the default sorting method.

-r
--reverse-sort

Sort output items in reverse order.

-s

--merge Merge the summarized accounting data into the summary files savacct and

          usracct.
-t
--print-ratio

For each entry, print the ratio of real time to the sum of system and user times. If the sum of system and user times is too small to report—the sum is zero—*ignore* will appear in this field.

-u
--print-users

For each command in the accounting file, print the userid and command name. After printing all entries, quit. Note: this flag supersedes all others.

-v num
--threshold num

Print commands which were executed num times or fewer and await a reply from the terminal. If the response begins with y, add the command to the

          **junk** group.
--separate-forks

It really doesn’t make any sense to me that the stock version of sa separates statistics for a particular executable depending on whether or not that command forked. Therefore, GNU sa lumps this information together unless this option is specified.

--sort-real-time

Sort the output by the “real time” (elapsed time) for each command.

--ahz hz Use this flag to tell the program what AHZ should be (in hertz). This option is useful if you are trying to view an acct file created on another machine which has the same byte order and file format as your current machine, but has a different value for AHZ.

--debug Print verbose internal information.

-V
--version

Print sa’s version number.

-h

--help Print sa’s usage string and default locations of system files to standard output.

Note: if more than one sorting option is specified, the list will be sorted by the one specified last on the command line.

5.2 Problems

I haven’t been able to test this on many different machines because the data files grow so big in a short time; our sysadmin would rather save the disk space.

Most versions of sa that I’ve tested don’t pay attention to flags like --print-seconds and --sort-num-calls when printing out commands when combined with the --user-summary or --print-users flags. GNU sa pays attention to these flags if they are applicable.

5.2.1 mips sa

The average memory use is stored as a short rather than a double, so we suffer from round-off errors. GNU sa uses double the whole way through.

Table of Contents

Preface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

A Note on File Names and Locations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 History of the Accounting Utilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

1 ac. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

1.1 Flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.2 Problems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

The FTP Problem. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 The Shutdown/Reboot Problem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Stupid System V Machines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

2 accton. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

2.1 Flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

3 last. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

3.1 Flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.2 Problems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

The Clock Change Problem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 The Ftp Problem. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

4 lastcomm. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

4.1 Flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

5 sa . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

5.1 Flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 5.2 Problems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

5.2.1 mips sa . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17